OpenStudy (anonymous):

find the inverse Laplace transform of, 1/s2+a2

OpenStudy (anonymous):

is the equation (1/s^2) + a^2 or 1/(s^2 + a^2)??

OpenStudy (anonymous):

1/(s^2 + a^2)=(1/a) sin at ans,,,..

OpenStudy (anonymous):

hope that help dodos..lol

OpenStudy (anonymous):

inverse Laplace transform of \[1\over{s ^{2}+a ^{2}}\] is \[{1\over a} \sin(at)\]. mark o. gave you the correct answer